# Evolution of a Single Impurity Across the Superfluid–Mott insulator Transition in the Bose–Hubbard Model

### v1

@@ III Methods and observables
The algorithm introduces two independent worm operators that allow creation, moving, and annihilation of worldline discontinuities on each species, enabling efficient sampling of both diagonal and off-diagonal correlators. Because the impurity interacts only via density–density terms and both species are bosonic, the method is strictly sign-problem free for all parameters studied here.
@@
In the following, we summarize the observables used to characterize the impurity mobility and the accompanying bath deformation. All measurements are performed in the (canonical) single-impurity sector and in the grand-canonical ensemble for the bath, using the multi-species worm algorithm. Three global observables are central to our analysis. First, the impurity winding number $\langle W_{\mathrm{imp}}^{2}\rangle$ quantifies the coherent mobility of the impurity: a finite value signals a mobile light polaron, while its continuous suppression and eventual collapse indicate interaction-driven self-localization, ranging from heavy polarons and saturated bubbles or bound clusters in a superfluid bath to fully localized defects in a Mott-insulating bath. Second, the bath superfluid density $\rho_{\mathrm{s}}$, obtained from the standard winding-number estimator, tracks the loss of global phase stiffness and thus diagnoses the SF–MI transition and the underlying bath compressibility. Third, impurity-centered density responses, encoded in the correlator $C_{\mathrm{ib}}(\mathbf{r})$ and its cumulative bath-density deformation $\Delta N(R)$, resolve the real-space structure of the dressing cloud and distinguish depletion bubble, bound clusters, nearly free defect, and localized defect state. Statistical errors are estimated via jackknife analysis over independent Monte Carlo blocks, and radial profiles are smoothed by binning over shells of constant distance using the minimum-image convention.
@@ III.1 Impurity weight and bath baseline
and monitor $\langle W_{\mathrm{imp}}^{2}\rangle$ to distinguish between mobile and localized impurity regimes.
@@
$\kappa_{\mathrm{b}}=\frac{\beta}{L^{2}}\left(\langle N_{\mathrm{b}}^{2}\rangle-\langle N_{\mathrm{b}}\rangle^{2}\right),\qquad N_{\mathrm{b}}=\sum_{\mathbf{r}}\langle n_{\mathrm{b}}(\mathbf{r})\rangle.$ (6)
@@
$G_{\mathrm{imp}}(\mathbf{r},\tau)=\bigl\langle a_{\mathrm{imp}}(\mathbf{r},\tau)\,a_{\mathrm{imp}}^{\dagger}(\mathbf{0},0)\bigr\rangle,$ (7)
@@ III.2 Impurity-centered correlator and radial averaging
Here $\mathbf{r}_{\mathrm{imp}}$ labels the lattice site occupied by the impurity, $C_{\mathrm{ib}}(\mathbf{r})$ measures the change of bath density conditioned on the impurity location: $C_{\mathrm{ib}}(\mathbf{r})/t<0$ corresponds to depletion (repulsive $U_{\mathrm{ib}}/t>0$), while $C_{\mathrm{ib}}(\mathbf{r})/t>0$ corresponds to accumulation (attractive $U_{\mathrm{ib}}/t<0$).
@@ III.3 Cumulative excess and its characteristic measures
By construction $\Delta N(R\!\to\!\infty)=0$ in the canonical ensemble (global particle number is fixed), while the behavior at small $R$ reveals the local deformation created by the impurity.
@@ III.4 On-site contrast
Together, $\langle W_{\mathrm{imp}}^{2}\rangle$, $C_{\mathrm{ib}}(R)$, $\Delta N(R)$, and the extremal measures $\{R(\Delta N_{\rm ext}),\,\Delta N_{\rm ext}\}$ provide a detailed microscopic characterization of impurity dressing and localization in both the SF and MI phases and across the SF–MI transition.
IV Attractive impurity in a superfluid bath: real-space self-localization
In our companion Letter [28], we showed that, in a compressible superfluid bath at fixed $U_{\mathrm{b}}/t=13.3$, impurity localization is governed by an interaction-driven winding-collapse mechanism: as a repulsive impurity–bath coupling $U_{\mathrm{ib}}/t>0$ is increased, the impurity evolves from a mobile light polaron with finite winding and an extended dressing cloud to a heavy self-trapped state and, eventually, to a depletion saturated-bubble configuration with $\Delta N(R)\to-1$. In this section we turn to the attractive side. To reveal the mirror counterpart of the repulsive depleted bubble and to resolve the real-space structure of self-localization, we again fix the bath interaction at $U_{\mathrm{b}}/t=13.3$ (in the superfluid regime) and tune the attractive impurity–bath coupling $U_{\mathrm{ib}}/t<0$.
Figure 2: Quasiparticle properties of an attractive impurity in a superfluid bath. (a) the polaron ground-state energy $E_{\mathrm{p}}/t$, (b) the effective-mass ratio $m^{*}/m_{0}$ with $m_{0}$ the bare particle mass, and (c) the quasiparticle residue $Z_{0}$ as functions of the attractive impurity–bath coupling $|U_{\mathrm{ib}}|/t$ at fixed bath interaction $U_{\mathrm{b}}/t=13.3$, deep in the superfluid regime. Error bars denote statistical uncertainties and are smaller than the symbols where not visible.
@@ IV Attractive impurity in a superfluid bath: real-space self-localization
We start with the interaction-driven path in the superfluid bath and track how the impurity quasiparticle evolves as the attractive impurity-bath coupling is increased at fixed $U_{\mathrm{b}}/t=13.3$. Figure 2 shows the polaron energy $E_{\mathbf{p}}/t$, the effective-mass ratio $m^{*}/m_{0}$, and the residue $Z$ as functions of $U_{\mathrm{ib}}/t<0$. For weak attraction $|U_{\mathrm{ib}}|/t<10.0$ the impurity forms a moderately dressed light polaron: the energy is only linearly shifted, the mass remains close to the bare value ($m^{*}/m_{0}\approx 1$$1.2$), and the residue is high, $Z\approx 1.0$$0.6$. As $|U_{\mathrm{ib}}|/t$ is increased, the energy drops almost linearly, the effective mass grows steadily to $m^{*}/m_{0}\gtrsim 1.5$, and the residue is reduced to $Z\sim 0.2$, indicating progressively stronger dressing. Upon entering the strong-attraction regime $|U_{\mathrm{ib}}|/t>13.3$, the mass shows a sharp upturn while the residue is strongly suppressed, signaling the onset of an interaction-driven self-trapped cluster with very limited coherent mobility. This quasiparticle evolution is fully consistent with the winding-number collapse and the real-space contraction of the dressing cloud discussed in Figure 3 to 5, and together they establish a continuous crossover from a mobile light polaron at weak attractive impurity-bath coupling $|U_{\mathrm{ib}}|/t$ to a heavy, nearly localized bound cluster at strong attraction in a still globally superfluid bath.
Figure 3: Real-space impurity dressing for attractive impurity in a superfluid bath at $U_{\mathrm{b}}/t=13.3$. (a) Cumulative bath-density enhancement $\Delta N(R)$ for various attractive impurity–bath couplings $U_{\mathrm{ib}}/t<0$, showing the evolution from a broad, weak dressing cloud (light polaron) to a sharply localized multi-particle cluster. (b) Radius of the maximum $R(\Delta N_{\max})$, demonstrating the monotonic inward compression of the dressing cloud as $|U_{\mathrm{ib}}|/t$ increases. (c) Magnitude of maximum $\Delta N_{\max}(R)$, which increases without saturation, reflecting the grand-canonical nature of the bath: stronger attraction continuously recruits additional bosons into the impurity vicinity. Unlike the repulsive saturated-bubble case where $\Delta N(R)$ saturates at 1-1, the attractive branch exhibits unbounded growth due to unrestricted local accumulation. These results demonstrate a continuous crossover: light polaron $\rightarrow$ heavy polaron $\rightarrow$ bound cluster, while the bath remains globally superfluid.
@@
For stronger attractive impurity–bath coupling ($|U_{\mathrm{ib}}|/t>20.0$), the deformation cloud collapses toward the impurity site and $\Delta N(R)$ develops a sharply peaked profile at $R\simeq 2$, indicating the formation of a tightly bound multi-particle cluster. Unlike the repulsive saturated-bubble case, where the cumulative deficit $\Delta N(R)$ rapidly saturates at approximately 1-1 once $R$ exceeds the depletion radius, here the cumulative excess $\Delta N_{\max}$ continues to grow with $|U_{\mathrm{ib}}|/t$. Both branches are simulated in the same grand-canonical ensemble for the bath; the qualitative difference originates from the energetics of adding versus removing particles. For a repulsive impurity it is energetically favorable to expel essentially one bath boson from the impurity vicinity, producing a saturated depletion bubble with $\Delta N_{\min}\simeq-1$. For attractive $U_{\mathrm{ib}}/t<0$, the local potential well can instead bind several bath bosons: as $|U_{\mathrm{ib}}|/t$ increases the optimal on-site occupation near the impurity grows to $C(R=0)\approx 3\text{--}4$ for our strongest couplings, so the bound cluster recruits a few extra particles from the reservoir. In a $20\times 20$ system this corresponds to a total bath occupation of $N_{\mathrm{b}}\simeq 403$, i.e., a global filling only slightly above unity, while the local excess is fully captured by the cumulative bath-density response $\Delta N(R)$. Taken together, these results establish the continuous crossover within the superfluid bath, $\textit{light polaron}\;\rightarrow\;\textit{heavy polaron}\;\rightarrow\;\textit{bound cluster}$, all occurring while the bath remains globally superfluid.
@@
Figure 4: Real-space impurity-centered correlator $C_{\mathrm{ib}}(R)$ in the superfluid bath at $U_{\mathrm{b}}/t=13.3$ for various attractive couplings $U_{\mathrm{ib}}/t<0$. For weak attraction ($|U_{\mathrm{ib}}|/t\lesssim 13.3$, red curves), $C_{\mathrm{ib}}(R)$ is small and spatially broad, indicating an extended, compressible dressing cloud characteristic of a mobile light polaron. In the intermediate regime ($13.3<|U_{\mathrm{ib}}|/t\lesssim 20.0$, yellow curves), the response becomes stronger and short-ranged, signaling a compact heavy polaron with reduced coherent motion. For strong attraction ($|U_{\mathrm{ib}}|/t>20.0$, blue curves), the correlation collapses to a sharply peaked on-site form, forming a self-trapped, multi-particle bound cluster. The inset shows that the on-site value $C_{\mathrm{ib}}(0)$ grows monotonically with $|U_{\mathrm{ib}}|/t$ and does not saturate, highlighting that in a grand-canonical superfluid bath, a deeper attractive impurity continuously recruits additional bath bosons without intrinsic upper bound. This demonstrates an interaction-driven winding-collapse crossover in a compressible superfluid host.
@@
In Fig. 4, we examine the impurity-centered correlation function $C_{\mathrm{ib}}(R)$, which provides real-space insight into how the impurity distorts its local environment. The bath interaction is fixed at $U_{\mathrm{b}}/t=13.3$, well inside the superfluid phase, while the attractive impurity–bath coupling $U_{\mathrm{ib}}/t<0$ is tuned from weak to strong. For weak attraction ($|U_{\mathrm{ib}}|/t\lesssim 13.3$, red curves), $C_{\mathrm{ib}}(R)$ displays a shallow and spatially broad profile, reflecting a compressibility-enabled, extended dressing cloud characteristic of a coherent light polaron with finite winding and significant mobility, consistent with Fig. 5(a). In the intermediate regime ($13.3<|U_{\mathrm{ib}}|/t\lesssim 20$, yellow curves), the correlation grows in amplitude while the winding has already collapsed to zero, signaling the formation of a compact heavy-polaron state with strongly reduced coherent motion. For strong attraction ($|U_{\mathrm{ib}}|/t\gtrsim 20$, blue curves), all curves converge to a sharply peaked, nearly universal shape with a dominant on-site peak and a pronounced nearest-neighbor maximum, indicative of a tightly bound multi-particle cluster. The inset shows that the on-site signal $C_{\mathrm{ib}}(0)$ increases monotonically with $|U_{\mathrm{ib}}|/t$ and does not saturate. This behavior reflects the grand-canonical nature of the superfluid bath at fixed chemical potential: a deeper attractive potential can continuously recruit additional bath bosons into the cluster. It contrasts sharply with the repulsive branch, where the cumulative bath-density response $\Delta N(R)$ saturates at approximately 1-1 once a single bath particle has been expelled from the impurity vicinity.
@@
Taken together, these results demonstrate a continuous, interaction-driven winding-collapse crossover from a coherent light polaron to a heavy polaron and finally a self-trapped, multi-particle bound cluster, even though the bath remains globally superfluid. The collapse of the spatial dressing profile thus provides a real-space signature of interaction-driven localization at fixed bath compressibility.
@@
Figure 5: Interaction-controlled winding-collapse localization of an attractive impurity in a superfluid bath at fixed $U_{\mathrm{b}}/t=13.3$. (a) Impurity winding number $\langle W_{\mathrm{imp}}^{2}\rangle$ as a function of attractive coupling $U_{\mathrm{ib}}/t<0$. A smooth suppression followed by a rapid collapse at $|U_{\mathrm{ib}}|/t\gtrsim 10$ signals a transition from a mobile light polaron to a localized bound cluster. (b) Bath superfluid density $\rho_{\mathrm{b}}$ remains essentially unchanged across the entire coupling range, confirming that the impurity localization is not induced by the loss of global coherence in the bath. This establishes an interaction-driven winding-collapse localization mechanism: the impurity becomes localized even while the bath remains fully superfluid and compressible.
@@
Winding-controlled localization of an attractive impurity in the superfluid bath.
@@
Figure 5(a) shows that $\langle W_{\mathrm{imp}}^{2}\rangle$ decreases smoothly at first, then rapidly collapses by nearly two orders of magnitude once $|U_{\mathrm{ib}}|/t\gtrsim 13.3$, signaling a crossover from a mobile light polaron to a heavy polaron. This loss of winding takes place without any qualitative change in the bath, as confirmed in Fig. 5 (b), where the bath superfluid density $\rho_{\mathrm{b}}$ remains essentially constant across the entire range of impurity-bath couplings. The coexistence of vanishing impurity winding with a finite bath superfluid density clearly demonstrates that localization is not driven by the breakdown of global phase coherence in the bath, but rather by an interaction-driven winding collapse, i.e., self-trapping of the impurity.
In contrast to a Mott-insulating background, where impurity localization involves quantized defect formation and abrupt expulsion or recruitment of a single bath boson, here the collapse of mobility is continuous. The bath remains fully superfluid and compressible, yet the impurity loses all long-range coherence and becomes localized purely by increasing the depth of the local attractive impurity-bath coupling. This establishes a clear case of interaction-driven winding-collapse localization in a superfluid bath: the impurity loses its ability to form worldline windings even though the bath retains macroscopic superfluidity.

### v2

@@ III Methods and observables
The algorithm introduces two independent worm operators that allow creation, moving, and annihilation of worldline discontinuities on each species, enabling efficient sampling of both diagonal and off-diagonal correlators.
@@
In the following, we summarize the observables used to characterize the impurity mobility and the accompanying bath deformation. All measurements are performed in the (canonical) single-impurity sector and in the grand-canonical ensemble for the bath, using the multi-species worm algorithm. Three global observables are central to our analysis. First, the impurity winding number $\langle W_{\mathrm{imp}}^{2}\rangle$ quantifies the coherent mobility of the impurity: a finite value signals a mobile light polaron, while its continuous suppression and eventual collapse indicate interaction-driven self-trapping, ranging from heavy polarons and saturated bubbles or bound clusters in a superfluid bath to fully self-trapped defects in a Mott-insulating bath. Second, the bath superfluid density $\rho_{\mathrm{b}}$, obtained from the standard winding-number estimator, tracks the loss of global phase stiffness and thus diagnoses the SF–MI transition. Third, impurity-centered density responses, encoded in the correlator $C_{\mathrm{ib}}(R)$ and its cumulative bath-density deformation $\Delta N(R)$, resolve the real-space structure of the dressing cloud and help distinguish depletion bubbles, bound clusters, nearly free defects, and self-trapped (defect-bound) states. Statistical errors are estimated via jackknife analysis over independent Monte Carlo blocks.
@@ III.1 Impurity weight and bath baseline
and monitor $\langle W_{\mathrm{imp}}^{2}\rangle$ to distinguish between mobile and self-trapped impurity regimes.
@@
$\kappa_{\mathrm{b}}=\frac{\beta}{L^{2}}\left(\langle N_{\mathrm{b}}^{2}\rangle-\langle N_{\mathrm{b}}\rangle^{2}\right),\qquad N_{\mathrm{b}}=\sum_{\mathbf{r}}n_{\mathrm{b}}(\mathbf{r}).$ (6)
@@
$G_{\mathrm{imp}}(\mathbf{r},\tau)=\bigl\langle a(\mathbf{r},\tau)\,a^{\dagger}(\mathbf{0},0)\bigr\rangle,$ (7)
@@ III.2 Impurity-centered correlator and radial averaging
Here $\mathbf{r}_{\mathrm{imp}}$ labels the lattice site occupied by the impurity, $C_{\mathrm{ib}}(\mathbf{r})$ measures the change of bath density conditioned on the impurity location: $C_{\mathrm{ib}}(\mathbf{r})<0$ corresponds to depletion (repulsive $U_{\mathrm{ib}}/t>0$), while $C_{\mathrm{ib}}(\mathbf{r})>0$ corresponds to accumulation (attractive $U_{\mathrm{ib}}/t<0$).
@@ III.3 Cumulative excess and its characteristic measures
By construction, $\Delta N(R\!\to\!\infty)=0$ (up to statistical noise), because $C_{\mathrm{ib}}(\mathbf{r})$ is defined relative to the global baseline $\bar{n}_{\mathrm{b}}$ within the same ensemble, while the behavior at small $R$ reveals the local deformation created by the impurity.
@@ III.4 On-site contrast
Together, $\langle W_{\mathrm{imp}}^{2}\rangle$, $C_{\mathrm{ib}}(R)$, $\Delta N(R)$, and the extremal measures $\{R(\Delta N_{\rm ext}),\,\Delta N_{\rm ext}\}$ provide a detailed microscopic characterization of impurity dressing and self-trapping in both the SF and MI phases and across the SF–MI transition.
IV Attractive impurity in a superfluid bath: real-space self-trapping
In our companion Letter [28], we showed that, in a compressible superfluid bath at fixed $U_{\mathrm{b}}/t=13.3$, impurity self-trapping is governed by an interaction-driven winding-collapse mechanism: as a repulsive impurity–bath coupling $U_{\mathrm{ib}}/t>0$ is increased, the impurity evolves from a mobile light polaron with finite winding and an extended dressing cloud to a heavy self-trapped state and, eventually, to a depletion saturated-bubble configuration with $\Delta N(R)\to-1$. In this section we turn to the attractive side. To reveal the mirror counterpart of the repulsive depleted bubble and to resolve the real-space structure of self-trapping, we again fix the bath interaction at $U_{\mathrm{b}}/t=13.3$ (in the superfluid regime) and tune the attractive impurity–bath coupling $U_{\mathrm{ib}}/t<0$.
Figure 2: Quasiparticle properties of an attractive impurity in a superfluid bath. (a) the polaron ground-state energy $E_{\mathrm{p}}/t$, (b) the effective-mass ratio $m^{*}/m_{0}$ with $m_{0}$ the bare lattice mass of the impurity, and (c) the quasiparticle residue $Z_{0}$ as functions of the attractive impurity–bath coupling $|U_{\mathrm{ib}}|/t$ at fixed bath interaction $U_{\mathrm{b}}/t=13.3$, deep in the superfluid regime. Error bars denote statistical uncertainties and are smaller than the symbols where not visible.
@@ IV Attractive impurity in a superfluid bath: real-space self-trapping
We start with the interaction-driven path in the superfluid bath and track how the impurity quasiparticle evolves as the attractive impurity-bath coupling is increased at fixed $U_{\mathrm{b}}/t=13.3$. Figure 2 shows the polaron energy $E_{\mathbf{p}}/t$, the effective-mass ratio $m^{*}/m_{0}$, and the residue $Z_{0}$ as functions of $U_{\mathrm{ib}}/t<0$. For weak attraction $|U_{\mathrm{ib}}|/t<10.0$ the impurity forms a moderately dressed light polaron: the energy is only linearly shifted, the mass remains close to the bare value ($m^{*}/m_{0}\approx 1$$1.2$), and the residue is high, $Z_{0}\approx 1.0$$0.6$. As $|U_{\mathrm{ib}}|/t$ is increased, the energy drops almost linearly, the effective mass grows steadily to $m^{*}/m_{0}\gtrsim 1.5$, and the residue is reduced to $Z_{0}\sim 0.2$, indicating progressively stronger dressing. Upon entering the strong-attraction regime $|U_{\mathrm{ib}}|/t>13.3$, the mass shows a sharp upturn while the residue is strongly suppressed, signaling the heavy polaron formation with very limited coherent mobility. This quasiparticle evolution is fully consistent with the winding-number collapse and the real-space contraction of the dressing cloud discussed in Figure 3 to 5, and together they establish a continuous crossover from a mobile light polaron at weak attractive impurity-bath coupling $|U_{\mathrm{ib}}|/t$ to a heavy polaron, finally a nearly self-trapped bound cluster at strong attraction in a still globally superfluid bath.
Figure 3: Real-space impurity dressing for attractive impurity in a superfluid bath at $U_{\mathrm{b}}/t=13.3$. (a) Cumulative bath-density enhancement $\Delta N(R)$ for various attractive impurity–bath couplings $U_{\mathrm{ib}}/t<0$, showing the evolution from a broad, weak dressing cloud (light polaron) to a sharply self-trapped multi-particle cluster. (b) Radius of the maximum $R(\Delta N_{\max})$, demonstrating the monotonic inward compression of the dressing cloud as $|U_{\mathrm{ib}}|/t$ increases. (c) Magnitude of maximum $\Delta N_{\max}(R)$, which increases without saturation, reflecting the grand-canonical nature of the bath: stronger attraction continuously recruits additional bosons into the impurity vicinity. Unlike the repulsive saturated-bubble case where $\Delta N(R)$ saturates at 1-1, the attractive branch exhibits unbounded growth due to unrestricted local accumulation. These results demonstrate a continuous crossover: light polaron $\rightarrow$ heavy polaron $\rightarrow$ bound cluster, while the bath remains globally superfluid.
@@
For stronger attractive impurity–bath coupling ($|U_{\mathrm{ib}}|/t>20.0$), the deformation cloud collapses toward the impurity site and $\Delta N(R)$ develops a sharply peaked profile at $R\simeq 2$, indicating the formation of a tightly bound multi-particle cluster. Unlike the repulsive saturated-bubble case, where the cumulative deficit $\Delta N(R)$ rapidly saturates at approximately 1-1 once $R$ exceeds the depletion radius, here the cumulative excess $\Delta N_{\max}$ continues to grow with $|U_{\mathrm{ib}}|/t$. Both branches are simulated in the same grand-canonical ensemble for the bath; the qualitative difference originates from the energetics of adding versus removing particles. For a repulsive impurity it is energetically favorable to expel essentially one bath boson from the impurity vicinity, producing a saturated depletion bubble with $\Delta N_{\min}\simeq-1$. For attractive $U_{\mathrm{ib}}/t<0$, the local potential well can instead bind several bath bosons: as $|U_{\mathrm{ib}}|/t$ increases the optimal on-site occupation near the impurity grows to $C_{\rm ib}(R=0)\approx 3\text{--}4$ for our strongest couplings, so the bound cluster recruits a few extra particles from the reservoir. In a $20\times 20$ system this corresponds to a total bath occupation of $N_{\mathrm{b}}\simeq 403$, i.e., a global filling only slightly above unity, while the local excess is fully captured by the cumulative bath-density response $\Delta N(R)$. Taken together, these results establish the continuous crossover within the superfluid bath, $\textit{light polaron}\;\rightarrow\;\textit{heavy polaron}\;\rightarrow\;\textit{bound cluster}$, all occurring while the bath remains globally superfluid.
@@
Figure 4: Real-space impurity-centered correlator $C_{\mathrm{ib}}(R)$ in the superfluid bath at $U_{\mathrm{b}}/t=13.3$ for various attractive couplings $U_{\mathrm{ib}}/t<0$. For weak attraction ($|U_{\mathrm{ib}}|/t\lesssim 13.3$, red curves), $C_{\mathrm{ib}}(R)$ is small and spatially broad, indicating an extended, compressible dressing cloud characteristic of a mobile light polaron. In the intermediate regime ($13.3<|U_{\mathrm{ib}}|/t\lesssim 20.0$, orange curves), the response becomes stronger and short-ranged, signaling a compact heavy polaron with reduced coherent motion. For strong attraction ($|U_{\mathrm{ib}}|/t>20.0$, blue curves), the correlation collapses to a sharply peaked on-site form, forming a self-trapped, multi-particle bound cluster. The inset shows that the on-site value $C_{\mathrm{ib}}(0)$ grows monotonically with $|U_{\mathrm{ib}}|/t$ and does not saturate, highlighting that in a grand-canonical superfluid bath, a deeper attractive impurity continuously recruits additional bath bosons without intrinsic upper bound. This demonstrates an interaction-driven winding-collapse crossover in a compressible superfluid host.
@@
In Fig. 4, we examine the impurity-centered correlation function $C_{\mathrm{ib}}(R)$, which provides real-space insight into how the impurity distorts its local environment. The bath interaction is fixed at $U_{\mathrm{b}}/t=13.3$, well inside the superfluid phase, while the attractive impurity–bath coupling $U_{\mathrm{ib}}/t<0$ is tuned from weak to strong. For weak attraction ($|U_{\mathrm{ib}}|/t\lesssim 13.3$, red curves), $C_{\mathrm{ib}}(R)$ displays a shallow and spatially broad profile, reflecting a compressibility-enabled, extended dressing cloud characteristic of a coherent light polaron with finite winding and significant mobility, consistent with Fig. 5(a). In the intermediate regime ($13.3<|U_{\mathrm{ib}}|/t\lesssim 20$, orange curves), the correlation grows in amplitude while the winding has already collapsed to zero, signaling the formation of a compact heavy-polaron state with strongly reduced coherent motion. For strong attraction ($|U_{\mathrm{ib}}|/t\gtrsim 20$, blue curves), all curves have a dominant on-site peak and a pronounced nearest-neighbor maximum, indicative of a tightly bound multi-particle cluster. The inset shows that the on-site signal $C_{\mathrm{ib}}(0)$ increases monotonically with $|U_{\mathrm{ib}}|/t$ and does not saturate. This behavior reflects the grand-canonical nature of the superfluid bath at fixed chemical potential: a deeper attractive potential can continuously recruit additional bath bosons into the cluster. It contrasts sharply with the repulsive branch, where the cumulative bath-density response $\Delta N(R)$ saturates at approximately 1-1 once a single bath particle has been expelled from the impurity vicinity.
@@
Taken together, these results demonstrate a continuous, interaction-driven winding-collapse crossover from a coherent light polaron to a heavy polaron and finally a self-trapped, multi-particle bound cluster, even though the bath remains globally superfluid. The collapse of the spatial dressing profile thus provides a real-space signature of interaction-driven self-trapping at fixed bath compressibility.
@@
Figure 5: Interaction-controlled winding-collapse self-trapping of an attractive impurity in a superfluid bath at fixed $U_{\mathrm{b}}/t=13.3$. (a) Impurity winding number $\langle W_{\mathrm{imp}}^{2}\rangle$ as a function of attractive coupling $U_{\mathrm{ib}}/t<0$. A smooth suppression followed by a rapid collapse at $|U_{\mathrm{ib}}|/t\gtrsim 10$ signals a transition from a mobile light polaron to a self-trapped bound cluster. (b) Bath superfluid density $\rho_{\mathrm{b}}$ remains essentially unchanged across the entire coupling range, confirming that the impurity self-trapping is not induced by the loss of global coherence in the bath. This establishes an interaction-driven winding-collapse self-trapping mechanism: the impurity becomes self-trapped even while the bath remains fully superfluid and compressible.
@@
Winding-controlled self-trapping of an attractive impurity in the superfluid bath.
@@
Figure 5(a) shows that $\langle W_{\mathrm{imp}}^{2}\rangle$ decreases smoothly at first, then rapidly collapses by nearly two orders of magnitude once $|U_{\mathrm{ib}}|/t\gtrsim 13.3$, signaling a crossover from a mobile light polaron to a heavy polaron. This loss of winding takes place without any qualitative change in the bath, as confirmed in Fig. 5 (b), where the bath superfluid density $\rho_{\mathrm{b}}$ remains essentially constant across the entire range of impurity-bath couplings. The coexistence of vanishing impurity winding with a finite bath superfluid density clearly demonstrates that self-trapping is not driven by the breakdown of global phase coherence in the bath, but rather by an interaction-driven winding collapse, i.e., self-trapping of the impurity.
In contrast to a Mott-insulating background, where impurity self-trapping involves quantized defect formation and abrupt expulsion or recruitment of a single bath boson, here the collapse of mobility is continuous. The bath remains fully superfluid and compressible, yet the impurity loses all long-range coherence and becomes self-trapped purely by increasing the depth of the local attractive impurity-bath coupling. This establishes a clear case of interaction-driven winding-collapse self-trapping in a superfluid bath: the impurity loses its ability to form worldline windings even though the bath retains macroscopic superfluidity.